# CLAP: Unsupervised 3D Representation Learning for Fusion 3D Perception via Curvature Sampling and Prototype Learning

### v2

@@ 4.2 Main Results
NuScenes Results. As shown in Table 1, CLAP achieves 2.48% mAP improvement over randomly initialization at convergence, which is $100\%$ more improvement for mAP than SOTA unsupervised 3D representation method UniPAD [58] and the best among all initialization methods. For NDS metric, UniPAD [58] only achieves comparable performance and PPKT [28] makes a gain of 0.57% while CLAP surpasses the train-from-scratch model by 1.76% . When it turns to different categories, CLAP generally benefit the performance of all the categories and for Construction Vehicle, Bus, Barrier, Motorcycle and Bicycle, the improvement over random initialization are more than 2% AP.
@@
Table 2: Results for fine-tuning on $1\%$ of training set in Waymo [45]. We first increase training iterations of from-scratch model and observe convergence (Rand.). Then we fix the training schedule to fine-tune the pre-training models. We provide an average difference $\bar{\Delta}$ to show the performance gain. All results are in %.
Waymo Results. In Table 2, we provide the average performance difference of mAP and mAPH over different difficulty levels. It can be found that CLAP achieves the best performance at convergence. Meanwhile, the performance gain brought by CLAP is approximately two times as the best (OCC-MAE [34]) of previous pre-training methods. This demonstrates the effectiveness and generalization ability of CLAP.
Potential Scaling Property. As we are not able to scale up the pre-training dataset at current stage, we explore potential scaling property by gradually decreasing the sample numbers (2.5%, 1% and 0.5%) for fine-tuning on NuScenes, which increases the ratio between pre-training data and fine-tuning data. The results are shown in Table 3. It can be found that as the ratio between pre-training data and fine-tuning data gets larger, the performance improvement by CLAP increases and CLAP provides a gain up to 7.22% mAP and 4.71% NDS with 0.5% fine-tuning data. These results show that CLAP is promising in scaling property and in the future, if we can scale up the pre-training dataset, CLAP might further improve current SOTA performance.
Init. Fine-tune Data mAP NDS
@@
Table 3: Results on potential scaling property of CLAP . “Fine-tune Data” stands for the sample number of training set in NuScenes [3]. Results for random initialization are at convergence and fine-tuning with CLAP uses the same numbers of iterations. Results are mAP and NDS in %.
Figure 4: Visualization of curvature estimation. The color of the points change from blue to red, indicating lower sample weight / curvature to higher ones. We highlight those correct regions with orange boxes and those noisy estimation with green ones. Some of the correct estimated regions are further zoom-in for better understanding. Best view in color.
4.3 Visualizations
We use the pre-trained model to estimate the curvature of LiDAR point clouds and use heatmap color to indicate the weight computed in Section 3.3. The visualization results are shown in Figure 4. We use orange boxes to highlight those regions with relatively correct estimation and green ones for those with noisy estimation. Furthermore, we zoom in some orange boxes. It can be found that though some noise exists (because pre-training is conducted in unsupervised manner), CLAP is able to predict high weights for those highly informative region for sampling and meanwhile assign lower weights to most of the background, which makes joint pre-training feasible.
@@ 4.3 Visualizations
4.4 Ablation Study
We conduct ablation study to evaluate the effectiveness of different components. Results are in Table 4. The first line is seperate pre-training with UniPAD [58]. The second one is jointly pre-training using UniPAD [58] and uniform sampling guided by depth (“Memory-friendly Ray Sampling” in [58]) to address the GPU memory limitation. It can be found that using simple sampling method does not bring improvement over separate pre-training. Then we add Curvature Sampling in third line and found that Curvature Sampling enables more effective sampling and improves the performance over uniform sampling guided by depth and separate pre-training. Finally, the Prototype Learning scheme (fourth line) learns a common feature space for segments of 3D scenes and introduces interaction of LiDAR and camera encoders, which achieves the best performance.
5 Conclusion
In this paper, we propose CLAP for unsupervised fusion joint pre-training via differentiable rendering. CLAP uses Curvature Sampling to sample more informative parts and learnable prototypes to represent parts of 3D scenes, optimized with an Expectation-Maximization approach. To further make use of the interplay between LiDAR geometry and image semantics, a swapping prediction loss is used. Finally, a Gram Matrix Regularization loss is proposed to avoid collapse. Experiment results demonstrate that CLAP is superior in unsupervised 3D representation learning and has the potential to scale up.

### v3

@@ 4.2 Main Results
NuScenes Results. As shown in Table 1, CLAP achieves 2.48% mAP improvement over randomly initialization at convergence, which is $100\%$ more improvement for mAP than SOTA unsupervised 3D representation method UniPAD (Yang et al., 2024) and the best among all initialization methods. For NDS metric, UniPAD (Yang et al., 2024) only achieves comparable performance and PPKT (Liu et al., 2021a) makes a gain of 0.57% while CLAP surpasses the train-from-scratch model by 1.76% . When it turns to different categories, CLAP generally benefit the performance of all the categories and for Construction Vehicle, Bus, Barrier, Motorcycle and Bicycle, the improvement over random initialization are more than 2% AP.
Waymo Results. In Table 4.2, we provide the average performance difference of mAP and mAPH over different difficulty levels. It can be found that CLAP achieves the best performance at convergence. Meanwhile, the performance gain brought by CLAP is approximately two times as the best (OCC-MAE (Min et al., 2023)) of previous pre-training methods. This demonstrates the effectiveness and generalization ability of CLAP.
Table 2: Results for Waymo (Sun et al., 2020).
@@
Table 3: Results on scaling property.
Init. F.T. Data mAP NDS
@@
Table 4: Ablation Study. The second line is joint pre-training with sampling method from UniPAD.
@@
Figure 4: Visualization of curvature estimation.
Potential Scaling Property. As we are not able to scale up the pre-training dataset at current stage, we explore potential scaling property by gradually decreasing the sample numbers (2.5%, 1% and 0.5%) for fine-tuning on NuScenes, which increases the ratio between pre-training data and fine-tuning data. The results are shown in Table 4.2 (F.T. is fine-tuning). It can be found that as the ratio between pre-training data and fine-tuning data gets larger, the performance improvement by CLAP increases and CLAP provides a gain up to 7.22% mAP and 4.71% NDS with 0.5% fine-tuning data. These results show that CLAP is promising in scaling property and in the future, if we can scale up the pre-training dataset, CLAP might further improve current SOTA performance.
4.3 Visualizations on Curvature Estimation
We use CLAP to estimate the curvature of point clouds and use heatmap color to indicate the weight computed in Section 3.3. As shown in Figure 4 (orange boxes for those regions with relatively correct estimation and green ones for those with noisy estimation), it can be found that though some noise exists (because pre-training is conducted in unsupervised manner), CLAP is able to predict high weights for those highly informative region for sampling and meanwhile assign lower weights to most of the background, which makes joint pre-training feasible.
4.4 Visualizations on Prototype Learning
We use the model pre-trained by CLAP to infer the 3D features and assign prototypes to different LiDAR points in the 3D space. Then we use different random colors to indicate different prototypes and visualize them, as shown in Figure 5. If can be found that the background road plane inside the same frame is generally assigned to the same prototype. And foreground vehicles are assigned to another prototype. This demonstrates that the proposed prototype learning scheme actually learns to represent parts of the scenes with prototypes in an unsupervised manner. However, as our pre-training does not incorporate any label, it can also be found that the prototype assignment has some noise, for example some of the road plane points are assigned to other prototypes.
Figure 5: Visualization of the prototype learning results. Different color indicates different prototype assignments.
4.5 Ablation Study
We conduct ablation study to evaluate the effectiveness of different components. Results are in Table 4.2. The first line is seperate pre-training with UniPAD (Yang et al., 2024). The second one is jointly pre-training using UniPAD (Yang et al., 2024) and uniform sampling guided by range (“Memory-friendly Ray Sampling” in (Yang et al., 2024)) to address the GPU memory limitation.
It can be found that using simple sampling method does not bring improvement over separate pre-training. Then we add Curvature Sampling in third line and found that Curvature Sampling enables more effective sampling and improves the performance over uniform sampling guided by range and separate pre-training. Finally, the Prototype Learning scheme (fourth line) learns a common feature space for segments of 3D scenes and introduces interaction of LiDAR and camera encoders, which achieves the best performance.
5 Conclusion
In this paper, we propose CLAP for unsupervised fusion joint pre-training via differentiable rendering. CLAP uses Curvature Sampling to sample more informative parts and learnable prototypes to represent parts of 3D scenes, optimized with an Expectation-Maximization approach. To explore the interplay between LiDAR geometry and image semantics, a swapping prediction loss is used with a regularization loss to avoid collapse. Experiment results demonstrate that CLAP is superior in unsupervised 3D representation learning and has the potential to scale up.
Acknowledgments
This paper is supported by the General Research Fund of Hong Kong No.17209324 and 17200622, and the Global Industrial Technology Cooperation Center (GITCC) through a grant agreement with the Korea Institute for Advancement of Technology (KIAT), project number P0028922.
Ethics statement
We further discuss potential negative social impact of CLAP in this section.
Job Displacement. Automation of tasks that require 3D perception, such as autonomous vehicles or robotics, might lead to job losses in sectors like transportation, warehousing, and manufacturing. While automation can create new jobs, the transition period can be challenging for those whose jobs are automated away.
Security Risks. The deployment of autonomous systems could introduce new vulnerabilities. Any failures in autonomous driving or robotics could lead to accidents or be exploited maliciously.
Accessibility and Inequality. The benefits of advanced 3D perception technologies might not be evenly distributed across society. Wealthier regions or organizations may have earlier and better access to these technologies, potentially widening the gap between different socioeconomic groups.
Reproducibility Statement
We provide method details in Section 4 and implementation details in Appendix A, which are enough for reproducing the experiments. Besides, we will release the code and models.
Appendix A More Implementation Details
Pre-training. We use a learning rate of 0.00005 with a cosine learning schedule for pre-training and use mask augmentation for CLAP with a masking rate of 0.9. All the pre-trainings are conducted on 8-H100 clusters for similar time ($\sim$ 45 mins / epoch). As pre-trained image backbones are broadly used for image feature extraction, the implementation of Yang et al. (2024) also use a pre-trained image backbone from Contributors (2020) to initiate their pre-training and the training set for this backbone does not involve any data from NuScenes Caesar et al. (2020) and Waymo Sun et al. (2020). We adopt this practice for training-from-scratch model and CLAP.
Downstream Training. We follow the common practice in OpenPCDet Team (2020) and only change the training iterations in order to observe the convergence of train-from-scratch models, which avoids the case that pre-training only accelerate convergence and make sure that pre-training indeed improve the performance of downstream models. The training epoch is 108 for BEVfusion Liu et al. (2023) with $5\%$ of NuScenes Caesar et al. (2020) training data and 252 for CenterPoint Yin et al. (2021) with $1\%$ of Waymo Sun et al. (2020) training data.
Appendix B More Experiment Results
B.1 Semantic Segmentation
@@
B.2 Repeated Evaluation
We use the same fixed random seed for all experiments in the main paper for reproducibility. As repeated evaluation can further reveal the training robustness, we repeat fine-tunings on NuScenes for 5 times with random initialization, UniPAD and CLAP. Mean and standard deviation of mAP are $48.55\pm 0.18\%$ (Rand.), $49.66\pm 0.29\%$ (UniPAD) and $51.16\pm 0.10\%$ (CLAP). CLAP achieves the best average performance and robustness against random seeds.
B.3 Transferring to Other Datasets
We conduct further experiments to evaluate the transferring ability of CLAP. Specifically, we select LiDAR-based 3D object detector CenterPoint Yin et al. (2021) on Once Mao et al. (2021) for the downstream task. A 40-beam LiDAR is utilized in Once Mao et al. (2021) to collect 15k labeled training data. We randomly sample $5\%$ and also use all of the labeled training set to train the from-scratch model until convergence is observed. Then we use pre-trained weights by CLAP on NuScenes Caesar et al. (2020) to initialize the same model and fine-tune it with the same training iterations as the randomly initialized model. Results are shown in Table 5. It can be found that pre-training by CLAP also benefits LiDAR-based 3D object detection, even in a cross-dataset setting. And if we look at the performance of “Rand*” and “CLAP*”, CLAP also accelerates the convergence in downstream task.
Init. F.T. mAP Vehicle Pedestrian Cyclist
0-30m 30-50m 50m- 0-30m 30-50m 50m- 0-30m 30-50m 50m-
Rand* 5% 20.48 58.03 25.22 12.98 11.62 9.75 6.97 21.55 6.83 3.11
CLAP* 22.86 +2.38 58.37 26.38 14.07 12.60 9.50 7.88 30.08 10.50 5.39
Rand 46.07 76.71 51.15 31.84 37.53 20.12 9.84 62.00 42.61 24.18
CLAP 46.88 +0.81 76.98 51.64 31.31 38.79 20.60 9.74 63.75 43.21 26.83
Rand* 100% 64.00 86.21 70.20 58.20 57.80 41.18 23.55 75.95 61.45 45.80
CLAP* 64.74 +0.74 88.14 72.59 59.13 57.37 42.24 24.22 77.11 61.91 45.63
Rand 65.03 88.18 74.23 61.75 57.32 38.90 21.96 78.07 64.32 48.16
CLAP 65.56 +0.53 87.97 72.77 62.11 58.33 40.11 21.29 78.63 64.70 47.27
Table 5: Results for transferring experiments on Once Mao et al. (2021) dataset. CenterPoint Yin et al. (2021) is used as the downstream detector. “Init.” indicates the initialization methods. “F.T.” indicates the number of training samples in fine-tuning stage. Overall mAP and APs for different categories within different ranges are shown in this table. “Rand*” means training the randomly initialized model with the original training iterations in OpenPCDet Team (2020). “Rand” indicates that we increase the number of training iterations for randomly initialized model until convergence is observed. “CLAP*” indicates that we pre-train the backbones with CLAP on NuScenes Caesar et al. (2020) and then fine-tune on Once with the original iterations in Team (2020). “CLAP” uses the same fine-tuning iterations as “Rand”. We use green color to highlight the performance improvement brought by CLAP. All the results are in %.
Init. mAP NDS mAP mAPH
Rand. 64.57 67.61 69.00 67.57
UniPAD 64.69 68.06 69.03 67.56
CLAP 65.11 68.38 69.49 68.04
Table 6: Fine-tuning on $100\%$ NuScenes and Waymo with fixed random seed.
B.4 More Fine-tuning Data
We fine-tune on $100\%$ NuScenes and Waymo with fixed random seed to provide more insights. Here, we show mAP and NDS for NuScenes in $2^{nd}$ and $3^{rd}$ cols and Level-2 mAP, mAPH for Waymo. CLAP improves random initialization by $\color[rgb]{0.1328125,0.546875,0.1328125}{\mathbf{+0.55}}$} whereas UniPAD by $\color[rgb]{0.1328125,0.546875,0.1328125}{\mathbf{+0.15}}$} on average, showing CLAP’s effectiveness. With Table 3 in the paper, more gain can be expected with more unlabeled data in real scenario.
Appendix C The Use of Large Language Models.
We use Large Language Models to help and aid editing/polishing the paper. In details, we polish the Introduction and Related Work sections with Large Language Models, which mostly focuses on grammar, spelling and word choice.